find the measure of an angle whose supplement is 45 degrees less than 4 times its complement.
Answer by mananth(12270)

Let the angle be x
supplement of angle = 180-x
Complement of angle = 90-x
4 times= 4(90-x)
4(90-x)-(180-x)=45
360-4x-180+x=45
135=3x
/3
x= 45
measure of angle = 45 deg